OpenStudy (anonymous):

7.What is the solution to the equation 9(x + 2) = 27 ? :p x = 2.5 x = 0.5 x = -0.5 x = -3.5

OpenStudy (callisto):

9(x + 2) = 27 1. divide both sides by 9 9(x + 2) = 27 9(x + 2)/9 = 27/9 x+2 = 3 2. substract 2 from both sides x+2-2 = 3 - 2 What do you get??

OpenStudy (anonymous):

That is stupid. The answer is 1, and it's not in the choices?

OpenStudy (anonymous):

9*3=27 :P

OpenStudy (anonymous):

9(1+3)=27

OpenStudy (callisto):

Any typo in the question? @skate2400

OpenStudy (anonymous):

1+2** my bad.

OpenStudy (anonymous):

no typo

OpenStudy (anonymous):



OpenStudy (anonymous):

9x = 18=27....27-18=9...9x=9...x=1.

OpenStudy (anonymous):

Must be typo :?

OpenStudy (callisto):

No typo => none of the above answers.

OpenStudy (anonymous):

loool.

OpenStudy (anonymous):

haha ur right it is my teacher's an idiot lol

OpenStudy (callisto):

Hmm... just a little mistake.... not an idiot :)